Find the solution of the difference equation y_(k+2)-4y_(k+1)+4y_k=0;k=0,1,2,...
Also find a particular solution when y_0=1 and y_1=6.
1
Expert's answer
2019-03-27T06:32:13-0400

(i) The characteristic equation is "C^2-4C+4=0".

The roots are "C_{1,2}=2\u2254C".

The general solution of the difference equation is "y_k=A_1C^k+A_2kC^k=A_12^k+A_2k2^k".


(ii) From the boundary conditions, we get that "A_1=1", "2A_1+2A_2=6".


Answer:

(i) "y_k=A_12^k+A_2k2^k".

(ii) "A_1=1", "A_2=2".


References:

1 "\url{https://en.wikipedia.org/wiki/Recurrence_relation#Solving_homogeneous_linear_recurrence_relations_with_constant_coefficients}"